# Navigating Data Corruption in Machine Learning: Balancing Quality, Quantity, and Imputation Strategies

### v1

@@ Abstract
Data corruption, including missing and noisy data, is an inevitable challenge in real-world machine learning applications. This paper investigates the impact of data corruption on model performance and explores strategies to mitigate its effects through two distinct experimental setups: supervised learning with NLP tasks (NLP-SL) and deep reinforcement learning for traffic signal optimization (Signal-RL). We evaluate the relationship between data corruption levels and model performance, assess the effectiveness of various data imputation methods, and analyze the utility of enlarging datasets to offset data corruption.
Our findings reveal that model performance under data corruption follows a diminishing return curve, well-modeled by the exponential function $S=a(1-e^{-b(1-p)})$, where $p$ is the corruption ratio. Missing data, while harmful, is less detrimental than noisy data, which causes rapid performance degradation and increased training instability, particularly in sequential decision-making tasks like Signal-RL. Imputation strategies introduce a trade-off between recovering missing information and introducing noise, with their effectiveness heavily dependent on imputation accuracy and corruption ratio. We identify a distinct “imputation advantageous corner” and “imputation disadvantageous edge” on an imputation advantage heatmap, and categorize tasks as either “noise-sensitive” or “noise-insensitive” based on their imputation decision boundaries.
Additionally, we show that increasing the dataset size mitigates, but does not fully eliminate, the effects of data corruption, especially under high noise levels. The marginal utility of additional data diminishes as corruption levels increase, confirming the exponential nature of the trade-off between data quality and quantity. An empirical rule emerges from our study: approximately 30% of the data is critical for determining model performance, while the remaining 70% has a minimal impact.
These findings provide practical insights into data preprocessing, imputation strategy selection, and resource allocation for data collection, offering actionable guidance for building robust machine learning systems in noisy, real-world environments.
@@ 1 Introduction
The advent of large language models (LLMs), such as GPT-3 and BERT, has revolutionized natural language processing (NLP), enabling machines to understand and generate human-like text with remarkable proficiency. Pre-training these models requires vast amounts of data to capture the nuances and complexities of human language. However, real-world datasets are often riddled with corrupted data, which can significantly impede the performance of LLMs if not properly addressed. Handling corrupted data is, therefore, a critical step in the pre-training pipeline, necessitating robust methods to ensure data integrity and model efficacy.
Corrupted data can arise from various sources, such as transmission errors or incomplete data collection processes. Data corruption manifests in different forms, including missing data and noisy data. The presence of corrupted data poses serious challenges to a model’s learning ability. Consequently, employing effective strategies to handle corrupted data is essential to maximize the potential of LLMs.
While numerous data imputation methods exist, their effectiveness in enhancing neural network learning efficiency remains unclear. LLMs have the inherent capability to learn directly from corrupted data, albeit at reduced efficiency. This raises important questions: What is the relationship between data corruption percentage and model performance? Can the “brute force” approach of adding more data compensate for poor data quality? If so, how much additional effort is required?
In this work, we explore the behavior of deep learning models under varying levels and types of data corruption. We conduct experiments on two types of tasks: supervised learning (NLP-SL) and reinforcement learning (Signal-RL). Additionally, we study the effectiveness of data imputation methods by comparing them against a no-imputation baseline. This study seeks to answer whether enlarging the dataset can mitigate the degradation caused by noise and, if so, how much additional data is needed. The insights gained from this study aim to provide guidance for data collection and preprocessing strategies.
The key takeaways from this paper may be summarized as follows:
  1. 1.
    Diminishing Returns in Data Quality Improvement: mproving data quality has diminishing returns on model performance. The relationship between the model score S and noise level p is well-modeled by the function:
    $S=a(1-e^{-b(1-p)})$
    where parameter $a=\frac{S_{0}}{1-e^{-b}}$; and $S_{0}$ is the model score when $p$ = 0 (i.e., no corruption).
  2. 2.
    Data Noise is More Detrimental than Missing Data: Noise causes more severe performance degradation compared to missing data. It also contributes to training instability, making the learning process less reliable.
  3. 3.
    Trade-offs in Data Imputation: Imputing missing data can introduce noise, resulting in a trade-off. The imputation advantage heatmap highlights two key regions:
  4. 4.
    • Noise-insensitive tasks: These tasks exhibit gradual performance degradation and are characterized by a decision boundary on the heatmap fitted by a sigmoid curve.
    • Noise-sensitive tasks: These tasks show sharp performance declines, with decision boundaries well-modeled by an exponential curve.
  5. 5.
    Limits of Enlarging Datasets: Data corruption leads to performance declines that cannot be fully recovered by simply increasing the sample size. Enlarging datasets does not entirely offset the detrimental effects of corruption.
  6. 6.
    Impact of Noise on Learning Efficiency: Data noise hampers learning efficiency. To achieve the same performance level (if at all possible), the number of required samples—and hence training time—increases exponentially with the noise level.
  7. 7.
    Empirical Rule on Data Importance: Approximately 30% of the data is critical for determining model performance, while the remaining 70% can be lost with minimal impact on performance. This finding provides practical guidance for data collection and prioritization.
@@ 2 Related Work
Types of Data Corruption
Data corruption in machine learning encompasses various forms, including missing data, noisy data, and adversarial perturbations. This study focuses on two major types of corruption: missing data and noisy data, which commonly occur in real-world scenarios.
Missing Data: Missing data can result from sensor dropout, incomplete data collection, or non-response in surveys. Rubin’s classification categorizes missing data into three types: Missing Completely at Random (MCAR), Missing at Random (MAR), and Missing Not at Random (MNAR) [Rub76]. In NLP, missing data often manifests as masked or unknown tokens, while in reinforcement learning, it appears as incomplete state observations.
Noisy Data: Data noise can affect both labels and features, with sources ranging from environmental factors to measurement errors. Noise is often categorized by its statistical distribution (e.g., Gaussian, uniform) or type (e.g., additive, multiplicative). Noisy data significantly disrupts model learning, particularly when noise affects key features or labels critical to decision-making.
Handling Corrupted Data
Data imputation aims to recover missing information, and various strategies have been proposed: • Statistical Methods: These include simple approaches like mean, mode, or median imputation [SG02], as well as more sophisticated techniques like regression imputation and multiple imputation [Rub87]. While straightforward, these methods may introduce bias or underestimate variability. • Machine Learning Methods: Techniques such as K-Nearest Neighbors (KNN) imputation [TCS+01], decision tree-based imputation [BFOS84], and Random Forest imputation [Bre01] leverage patterns in data to predict missing values. • Deep Learning-Based Imputation: Autoencoders and Generative Adversarial Networks (GANs) have emerged as powerful tools for data imputation. Denoising autoencoders reconstruct inputs with missing values [VLBM08], while GANs generate plausible synthetic data to fill gaps [GPAM+14]. Yuan et al. (2021) used masked language models (e.g., BERT) to impute missing tokens in text data [YWZ21].
For many applications, particularly LLM pre-training, masking missing data is often sufficient. Che et al. (2018) showed that masking missing time-series data combined with recurrent neural networks (e.g., GRU) can effectively handle missingness [CPC+18].
Impact of Data Corruption on Neural Network Learning
Supervised Learning (LLM Pre-training) Data corruption, including missing or noisy tokens, can significantly impair pre-trained language models like BERT and GPT. Missing data reduces the available context, weakening learned representations and hindering downstream tasks such as summarization or question answering [Dev18, Liu19]. Joshi et al. (2020) showed that missing rare tokens during pre-training leads to incomplete token embeddings, limiting the model’s ability to capture fine-grained semantics [JCL+20].
@@
Noisy data, particularly in large-scale corpora, introduces biases and degrades model robustness. Brown et al. (2020) highlighted that noisy training data increases the likelihood of generating biased or low-quality outputs, while filtering and robust training objectives mitigate such effects [Bro20]. Semantic noise, such as contradictory or irrelevant text, reduces the model’s ability to retain factual knowledge and generalize across tasks [PPF+20]. Studies by Gehman et al. (2020) further demonstrated that noisy datasets amplify harmful biases, underscoring the importance of pre-training data quality [GGS+20].
@@
Reinforcement Learning In reinforcement learning (RL), data corruption affects state observations, which are critical for decision-making. Missing features reduce state informativeness, leading to suboptimal policies, particularly in partially observable environments (POMDPs) [HS15]. Studies by Bai et al. (2021) demonstrated that missing features disrupt state transition dynamics, causing instability in model-free RL algorithms and inaccurate environment models in model-based RL [BGW19].
Noise in feature data further degrades RL performance. Pathak et al. (2017) showed that noisy features distort latent state representations, leading to poor decision-making in high-dimensional environments [PAED17]. Mnih et al. (2015) found that Q-values fluctuate with noisy observations, resulting in unstable policies [MKS+15]. Moreover, noise hampers transfer learning, as policies trained in corrupted environments fail to generalize to clean environments [TS09].
@@
  • Trade-Off Between Data Quality and Quantity: Can larger datasets compensate for data corruption? How much additional data is required to offset quality issues, and does the marginal utility of additional data diminish with increasing corruption?
This study aims to bridge the above gaps by evaluating the effects of data corruption on supervised and reinforcement learning tasks. We explore the utility of data imputation methods and analyze the trade-off between data quality and quantity, providing insights to guide data collection and preprocessing strategies.
@@ 3 Learning with corrupted data
Experiment Design
We designed two experiments, NLP supervised learning (NLP-SL) and traffic signal deep reinforcement learning (Signal-RL), to investigate the impact of data corruption. These two vastly different experimental setups were chosen because the questions we aim to explore are general and relevant across a wide range of machine learning tasks. By selecting tasks that are diverse in nature, we hope to derive more general insights and uncover deeper connections between these seemingly unrelated domains.
The first experiment focuses on the GLUE benchmark tasks. We use the pretrained BERT model as the base model and add feedforward layers on top to fine-tune it on eight tasks: CoLA, SST-2, MRPC, STSB, QQP, MNLI, QNLI, and RTE. The input data is corrupted by replacing certain words with the “[UNK]” token, while the labels $y$ remain uncorrupted. This type of data corruption is commonly encountered in natural language processing tasks. For instance, when digitizing text corpora, some words may become indistinguishable and are thus masked as unknown, whereas the associated labels are typically unaffected. Evaluating the amount of knowledge learned by a model remains a subjective challenge. In this experiment, we measure performance using classification-related accuracy metrics. Specifically:
  • The Matthews Correlation Coefficient (MCC) is used for CoLA.
  • The average of Pearson and Spearman correlation coefficients is used for STSB.
@@
Baseline scores (1/2 for binary classification and 1/3 for three-way classification) are subtracted from these metrics. The final model score is calculated as the average of the scores across all tasks. To ensure consistency, hyperparameters such as training epochs and learning rates are tuned using uncorrupted data and kept fixed for all subsequent experiments. Model convergence for these experiments is illustrated in Figure 2. For the sake of clarity, we will refer to this experiment as the “Natural Language Processing - Supervised Learning (NLP-SL)” experiment in the following sections.
The second experiment is a deep reinforcement learning (RL) task. We built an isolated intersection environment using SUMO. The environment is illustrated in Figure 1. The objective is to optimize the traffic signal at this intersection. The intersection consists of four approaches, each with three lanes. The traffic demand is generated using a binomial distribution, and the ratio of left-turn, through, and right-turn traffic demands is 1:3:2. The arrival rates are time-varying: East-West traffic demands follow a sine curve, while North-South demands follow a cosine curve, both within the interval $[0,\pi/2]$. The time horizon $H$ for the simulation is one hour.
A Deep Q-Network (DQN) model is used to learn the traffic signal control strategy. The state of the environment consists of road occupancy, the current signal phase, and the duration of the current signal phase, resulting in a state vector of dimension 965. A neural network with two hidden layers of sizes 256 and 48 is used to extract features. Layer normalization is applied, but no dropout layers are included. Standard techniques such as double networks and replay buffers are implemented. The action space is discrete, with four possible actions: East-West left turn, East-West through, North-South left turn, and North-South through. Each time step in the simulation corresponds to 6s. The reward $r_{t}$ for each time step is defined as the queuing vehicles’ value transformed according to Equation 1. The performance metric for the model is the episode cumulative reward, $R$. Although the reward is accumulated over one hour, the process itself is infinite.
@@
where $q_{t}$ is the number of queuing vehicles at time $t$; and the threshold for deciding whether a vehicle is stopped is $0.3m/s$.
We trained the model using linearly decaying exploration ($\epsilon$) and learning rate ($lr$) for 80% of the training time, after which the values were fixed. The initial and final $\epsilon$ values were 1.0 and 0.01, respectively, while the initial and final learning rates were 1e-3 and 1e-4, respectively. The DQN model convergence results are shown in Figure 2(b).
We experimented with three types of data corruption: vehicle-missing, inserting-noise, and masking-region:
  1. 1.
    Vehicle-missing: A proportion of the vehicles are not detected. This scenario is relevant in Vehicle-to-Everything (V2X) environments, where roadside units detect vehicles’ presence through communication channels like DSRC [THBM19]. However, only a proportion of vehicles are equipped with onboard devices. This type of corruption is analogous to the data missing scenario in the NLP-SL experiment.
  2. 2.
    Inserting-noise: Noise is added to the road occupancy state and rewards. This scenario is relevant in environments where vehicles are detected using computer vision systems, which can introduce errors.
  3. 3.
@@
Observations

### v2

@@ Abstract
Data corruption, including missing and noisy entries, is a common challenge in real-world machine learning. This paper examines its impact and mitigation strategies through two experimental setups: supervised NLP tasks (NLP-SL) and deep reinforcement learning for traffic signal control (Signal-RL). This study analyzes how varying corruption levels affect model performance, evaluate imputation strategies, and assess whether expanding datasets can counteract corruption effects. The results indicate that performance degradation follows a diminishing-return pattern, well modeled by an exponential function. Noisy data harms performance more than missing data, especially in sequential tasks like Signal-RL where errors may compound. Imputation helps recover missing data but can introduce noise, with its effectiveness depending on corruption severity and imputation accuracy. This study identifies clear boundaries between when imputation is beneficial versus harmful, and classify tasks as either noise-sensitive or noise-insensitive. Larger datasets reduce corruption effects but offer diminishing gains at high corruption levels. These insights guide the design of robust systems, emphasizing smart data collection, imputation decisions, and preprocessing strategies in noisy environments. The code for reproducing our experiments is available at https://github.com/qiliuchn/data-corruption-study.
@@ 1 Introduction
Machine learning models rely heavily on high-quality data, yet real-world datasets often suffer from corrupted data—such as missing entries or noisy measurements—which significantly degrades model performance. Data corruption arises from diverse sources, including sensor errors, transmission artifacts, or incomplete data collection. It manifests in two primary forms: missing data (e.g., masked tokens in NLP or undetected vehicles in traffic control) and noisy data (e.g., mislabeled text or perturbed sensor readings). While this challenge is well-documented in classical machine learning (Little and Rubin, 2019 [8]; Emmanuel et al., 2021 [4]), its implications for modern paradigms like large language models (LLMs) and deep reinforcement learning (DRL) remain underexplored.
Different learning paradigms exhibit distinct vulnerabilities to data corruption. For example, LLMs may tolerate certain missing tokens but struggle with semantic noise, whereas DRL policies can compound observational noise over sequential decisions. Prior work has shown that noise in NLP tasks (e.g., mislabeled text) biases model outputs (Brown et al., 2020 [31]), whereas in DRL, observational noise disrupts policy stability (Pathak et al., 2017 [39]). Despite these insights, fundamental questions remain unanswered: How do corruption type and severity impact different learning paradigms? Can imputation effectively recover lost performance, or does it can make things worse by introducing new noise? And when is collecting more data a viable alternative to cleaning existing data? While traditional imputation methods—including statistical interpolation and deep generative approaches—offer partial solutions, their efficacy varies significantly across tasks. This variation highlights persistent gaps in our understanding of the trade-offs between data quality, quantity, and imputation strategies in machine learning systems.
This study bridges these gaps through a systematic analysis of two distinct machine learning paradigms: supervised NLP tasks (NLP-SL) and deep reinforcement learning for traffic signal control (Signal-RL). It evaluates how varying corruption levels affect model performances, quantify the trade-offs of imputation strategies, and assess whether expanding datasets can counteract corruption effects. Experiments reveal universal patterns (such as diminishing returns from data quality improvements), paradigm-specific insights (such as DRL’s heightened sensitivity to noise), as well as task-specific observations (such as the critical 30% of data that determines performance in traffic signal control). This work provides actionable guidelines for designing robust systems in noisy environments, emphasizing smart data prioritization, imputation decisions, and noise-aware preprocessing. Key innovations include:
  • Modeling Performance Degradation: This study shows that performance degradation follows a diminishing-return pattern, well-captured by an exponential function, and reveal task-specific sensitivities.
  • Imputation Trade-offs: This study demonstrates the boundary conditions where imputation is beneficial versus harmful, providing actionable guidelines for practitioners.
  • Data Quantity-Quality Trade-offs: This study empirically shows that larger datasets can only partially offset corruption effects, with diminishing utility at high corruption levels.
By addressing these questions, our study advances the understanding of data corruption’s impact on machine learning and provides a framework for future research in robust model development.
@@ 2 Related Work
2.1 Types of Data Corruption
Data corruption in machine learning encompasses various forms, including missing data, noisy data, and adversarial perturbations. This study focuses on two major types of corruption: missing data and noisy data, which commonly occur in real-world scenarios. Missing data can result from sensor dropout, incomplete data collection, or non-response in surveys. Rubin’s classification categorizes missing data into three types: Missing Completely at Random (MCAR), Missing at Random (MAR), and Missing Not at Random (MNAR) [7]. In NLP, missing data often manifests as masked or unknown tokens, while in reinforcement learning, it appears as incomplete state observations. Data noise can affect both labels and features, with sources ranging from environmental factors to measurement errors. Noise is often categorized by its statistical distribution (e.g., Gaussian [2]/adversarial [29]) or types (e.g., additive/multiplicative [1]; label/feature [24]). Noisy data significantly disrupts model learning, particularly when noise affects key features or labels critical to decision-making.
2.2 Impacts of Data Corruption
@@
Prior research has empirically demonstrated the substantial impact of data corruption on model performance across learning paradigms. In the realm of language models, Joshi et al. (2020) [34] found that missing rare tokens during pre-training led to incomplete token embeddings that limits the model’s ability to capture fine-grained semantics. Missing data reduces the available context, weakening learned representations and hindering downstream tasks such as summarization or question answering [30, 42]. Noisy data, particularly in large-scale corpora, introduces biases and degrades model robustness. Brown et al. (2020) [31] highlighted that noisy training data increases the likelihood of generating biased or low-quality outputs, while filtering and robust training objectives mitigate such effects. Semantic noise, such as contradictory or irrelevant text, reduces the model’s ability to retain factual knowledge and generalize across tasks [41].
@@
For reinforcement learning, data corruption affects state observations, which are critical for decision-making. Missing features reduce state informativeness, leading to suboptimal policies, particularly in partially observable environments (POMDPs) [35]. Studies by Bai et al. (2019) [36] demonstrated that missing features disrupt state transition dynamics, causing instability in model-free RL algorithms and inaccurate environment models in model-based RL . Pathak et al. (2017) showed that noisy features distort latent state representations, leading to poor decision-making in high-dimensional environments [39]. Mnih et al. (2015) [38] found that Q-values fluctuate with noisy observations, resulting in unstable policies. Moreover, noise hampers transfer learning, as policies trained in corrupted environments fail to generalize to clean environments [40]. These studies collectively highlight how data corruption undermines performance in both paradigms, though through different mechanisms: contextual understanding and factual accuracy in language models versus compounding errors in sequential decision-making for reinforcement learning.
2.3 Data Imputation Techniques
Data imputation aims to recover missing information, and various strategies have been proposed. Emmanuel et al. [4] provide a comprehensive review of missing data in machine learning, classifying imputation methods into several categories: simple imputation, regression-based imputation, hot-deck imputation, the Expectation–Maximization (EM) method, multiple imputation, and machine learning–inspired techniques such as k-nearest neighbors (KNN), support vector machines (SVM), decision trees, clustering-based imputation, and ensemble methods. In a more recent taxonomy, Zhou et al. [3] group imputation approaches into four main types: statistical, machine learning–based, neural network–based, and optimization-based methods.
Statistical Methods: These include simple approaches like mean, mode, or median imputation [9], as well as more sophisticated techniques like regression imputation and multiple imputation [11]. While straightforward, these methods may introduce bias or underestimate variability.
Machine Learning Methods: Techniques such as K-Nearest Neighbors (KNN) imputation [14], decision tree-based imputation [15], and Random Forest imputation [16] leverage patterns in data to predict missing values. These methods often outperform simple statistical approaches by capturing complex nonlinear relationships in the data.
Deep Learning-based Imputation: Autoencoders and Generative Adversarial Networks (GANs) have emerged as powerful tools for data imputation. Denoising autoencoders reconstruct inputs with noisy values [17], while GANs generate plausible synthetic data to fill gaps [18]. Yuan et al. (2021) [21] used masked language models (e.g., BERT) to impute missing tokens in text data.
For many applications, particularly LLM pre-training, masking missing data is often sufficient. Che et al. (2018) showed that masking missing time-series data combined with recurrent neural networks (e.g., GRU) can effectively handle data missing [25]. We summarize common imputation methods, their strengths and weaknesses, and use cases in Table 1.
Table 1: Taxonomy of Imputation Techniques with Strengths, Weaknesses, and Use Cases
Category Method Strengths/Weaknesses Use Cases
Statistical-based [9, 11] Mean/Median/Mode + Simple, fast – Ignores correlations, distorts variance Small datasets, MCAR data
Maximum Likelihood + Handles MAR data well – Computationally intensive Surveys, clinical trials
Matrix Completion + Captures global structure – Requires low-rank assumption Recommendation systems
Bayesian Approach + Incorporates uncertainty – Needs prior distributions Small datasets with domain knowledge
Machine Learning [14, 15, 16] Regression-based + Models feature relationships – Assumes linearity Tabular data with correlations
KNN-based + Non-parametric, local patterns – Sensitive to $k$, scales poorly Small/medium datasets
Tree-based + Handles nonlinearity – Overfitting risk High-dimensional data
SVM-based + Robust to outliers – Kernel choice critical Nonlinear feature spaces
Clustering-based + Group-aware imputation – Depends on cluster quality Data with clear subgroups
Neural Network [17, 18, 21] ANN-based + Flexible architectures – Requires large data Complex feature interactions
Flow-based + Exact density estimation – High computational cost Generative tasks
VAE-based + Handles uncertainty – Blurry imputations Image/text incomplete data
GAN-based + High-fidelity samples – Training instability Media generation
Diffusion-based + State-of-the-art quality – Slow sampling High-stakes applications
@@ 2.3 Data Imputation Techniques
  • Trade-Off Between Data Quality and Quantity: Can larger datasets compensate for data corruption? How much additional data is required to offset quality issues and does the marginal utility of additional data diminish with increasing corruption level?
This study aims to bridge the above gaps by evaluating the effects of data corruption on supervised and reinforcement learning tasks. This study explores the utility of data imputation methods and analyze the trade-off between data quality and quantity, providing insights to guide data collection and preprocessing strategies.
@@ 3 Learning with corrupted data
3.1 Experiment Design
Two experiments are designed—Natural Language Processing Supervised Learning (NLP-SL) and Traffic Signal Deep Reinforcement Learning (Signal-RL)—to investigate the impact of data corruption. These two vastly different experimental setups were chosen to demonstrate the generality and broad relevance of the research questions across a wide range of machine learning tasks. By selecting tasks with diverse characteristics, we aim to derive more general insights and uncover deeper connections between these seemingly unrelated domains.
NLP Supervised Learning (NLP-SL)
The first experiment focuses on the GLUE benchmark tasks. Firstly, a BERT model is pre-trained using Wikitext and Bookcorpus as the base model; then the base model is frozen. We add classification head on top of the base model to fine-tune it on eight tasks: CoLA, SST-2, MRPC, STSB, QQP, MNLI, QNLI, and RTE using GLUE dataset. The GLUE input sentences are corrupted by replacing certain words with the [UNK] token, while the labels $y$ remain uncorrupted. This type of data corruption is commonly encountered in natural language processing tasks. For instance, when digitizing text corpora, some words may become indistinguishable and are thus masked as unknown, whereas the associated labels are typically unaffected. Evaluating the amount of knowledge learned by a model remains a subjective challenge. In this experiment, performance is measured using classification-related accuracy metrics. Specifically:
  • The Matthews Correlation Coefficient (MCC) is used for CoLA;
  • The average of Pearson and Spearman correlation coefficients is used for STSB;
@@ 3.1 Experiment Design
Baseline scores ($\frac{1}{2}$ for binary classification and $\frac{1}{3}$ for three-way classification) are subtracted from these metrics. The final model score is computed as the average of the scores across all tasks. To ensure consistency, hyperparameters such as the number of training epochs and learning rates are tuned using uncorrupted data and kept fixed for all subsequent experiments (see Table A1). Model convergence for these experiments is illustrated in Figure 2. For clarity, this experiment is referred to as the “Natural Language Processing—Supervised Learning (NLP-SL)” experiment in the following sections. Note that the base model is used only for sentence embedding and is frozen; the goal is to test the model’s ability to extract knowledge from corrupted data and store it in the feed-forward network.
For the NLP-SL task, two types of data corruption are introduced:
  • Data-missing: Each word in the training samples has a probability $p$ of being replaced with a [MASK] token.
  • Inserting-noise: Each word in the training samples has a probability $p$ of being replaced with a randomly selected word from the vocabulary.
Traffic Signal Control Deep Reinforcement Learning (Signal-RL)
The second experiment is a deep reinforcement learning (DRL) task. An isolated intersection environment is built using SUMO. The environment is illustrated in Figure 1. The objective is to optimize the traffic signal at this intersection. The intersection consists of four approaches, each with three lanes. The traffic demand is generated using a binomial distribution, and the ratio of left-turn, through, and right-turn traffic demands is 1:3:2. The arrival rates are time-varying: East-West traffic demands follow a sine curve, while North-South demands follow a cosine curve within range $[0,\pi/2]$. The simulation step size is one second, and each episode has a horizon $H$ of one hour.
A Deep Q-Network (DQN) model (Table A1) is used to learn the traffic signal control strategy. The state of the environment consists of road occupancy, the current signal phase, and the duration of the current signal phase, resulting in a state vector of dimension 965. A neural network with two hidden layers of sizes 256 and 48 is used to extract features. Layer normalization is applied, but no dropout layers are included. Standard techniques such as double networks and replay buffers are implemented. The action space is discrete, with four possible actions: East-West left turn, East-West through, North-South left turn, and North-South through. Each action in the simulation lasts for 6 seconds. The reward $r_{t}$ for each time step is defined as the number of queuing vehicles transformed according to Equation 1. The queue length is divided by 80 to normalize step rewards approximately within the range of 0 to 1. The performance indicator for the model is the episode cumulative reward, $R$. Although the reward is accumulated over one hour, the process itself is infinite.
@@
where $q_{t}$ is the number of queuing vehicles at the intersection at time $t$; and the threshold for deciding whether a vehicle is stopped is $0.3m/s$. $H$ is the horizon of one simulation episode.
The model is trained using linearly decaying exploration ($\epsilon$) and learning rate ($lr$) for 80% of the training time, after which the values were fixed. The initial and final $\epsilon$ values were 1.0 and 0.01, respectively, while the initial and final learning rates were 1e-3 and 1e-4, respectively. The DQN model convergence results are shown in Figure 2(b). The overview of model, dataset, and training configurations for both NLP-SL and Signal-RL tasks are shown by Table A1.
For Signal-RL task, three types of data corruption are introduced: vehicle-missing, inserting-noise, and masking-region:
  • Vehicle-missing: Each vehicle is not detected with probability $p$. This scenario is relevant in Vehicle-to-Everything (V2X) environments, where roadside units detect vehicles’ presence through communication channels like DSRC [43]. However, only a proportion of vehicles are equipped with onboard devices. This type of corruption is analogous to the data missing scenario in the NLP-SL experiment.
  • Inserting-noise: Noise is added to the road occupancy state and rewards. Each road cell occupancy state has probability $p$ of being replaced with random binary value. This scenario is relevant in environments where road occupancies are detected using computer vision systems, which can introduce errors.
@@
When investigating imputation methods, two common types of data-missing scenarios often encountered in practice are distinguished:
  • Exact imputation: This scenario arises when the precise locations of missing data are known. A common example occurs in natural language processing (NLP), where missing words are explicitly marked with placeholders such as “[UNK]”.
  • General imputation: In this case, the locations of missing data are unknown. For instance, in the Signal-RL experiment, vehicles may go undetected, leaving it unclear which elements of the state vector are corrupted. Imputing data under such conditions requires checking all possible locations, potentially introducing significantly more noise compared to exact imputation.
Figure 1: Simulation environment for Signal-RL experiment. The intersection comprises four approaches, each with three lanes. Traffic demand is generated based on a binomial distribution, with a left-turn:through:right-turn ratio of 1:3:2. The environment state is defined by road occupancy. Each simulation step corresponds to one second, and each episode spans one hour. The step-wise reward is defined in Equation 1.
Figure 2: Model convergence in two experiments. (a) NLP-SL experiment: The $x$-axis represents training steps, and the $y$-axis shows validation accuracy. (b) Signal-RL experiment: The $x$-axis denotes training episodes, and the $y$-axis indicates episode reward. Training configurations are detailed in Table A1.
3.2 Observations